# Semantic Agreement Enables Efficient Open-Ended LLM Cascades

### v1

@@ H.4 Performances at 40% Target Model FLOPs
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B
0.7337 0.734 0.7336 0.7341 0.7382 0.7343
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large
0.7345 0.7351 0.734 0.735 0.7402 0.7353
Qwen2.5-1.5B, Gemma3-1B,
FLAN-T5-Large
0.7366 0.7366 0.7359 0.737 0.7415 0.7376
Qwen2.5-1.5B, Llama3.2-1B,
mT0-Large
0.731 0.7323 0.7312 0.7328 0.7368 0.7335
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B, mT0-Large
0.7337 0.734 0.7338 0.7342 0.7398 0.7347
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large, FLAN-T5-Large
0.735 0.7361 0.734 0.7358 0.741 0.7359
Llama3.1-8B, Qwen2.5-7B,
FLAN-T5-Large
0.7443 0.7443 0.7431 0.744 0.7473 0.7449
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
0.7429 0.7431 0.7421 0.7432 0.7466 0.7446
Llama3.1-8B, Qwen2.5-7B,
Qwen2.5-1.5B
0.7438 0.7437 0.7443 0.744 0.7455 0.7444
Llama3.2-3B, Qwen2.5-3B,
Gemma3-1B
0.7379 0.7386 0.7377 0.7382 0.7423 0.7398
@@
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.684 0.6903 0.6887 0.1
Qwen2.5-3B 0.665 0.6745 0.673 0.2
Qwen2.5-1.5B 0.6579 0.6659 0.6667 0.3
Qwen2.5-0.5B 0.5947 0.6027 0.6014 0.3
Gemma3-1B 0.6664 0.6701 0.67 0.2
Llama3.2-1B 0.6467 0.6528 0.6522 0.4
Llama3.2-3B 0.6774 0.6872 0.6843 0.1
Llama3.1-8B 0.6909 0.6956 0.6968 0.6
FLAN-T5-Large 0.5834 0.5896 0.5894 0.4
mT0-Large 0.6107 0.6186 0.6154 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Gemma3-1B, Qwen2.5-1.5B 0.6673 0.6691 0.6705 0.1
Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6657 0.6692 0.6692 0.1
Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6625 0.6649 0.6663 0.0
Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.6645 0.6626 0.667 0.0
Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6592 0.6634 0.6638 0.1
Llama3.1-8B, Qwen2.5-7B 0.6869 0.6903 0.6898 0.1
Llama3.1-8B, Qwen2.5-3B 0.6818 0.6854 0.6876 0.2
Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6855 0.689 0.6892 0.1
Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6843 0.6883 0.6899 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B
0.6623 0.6622 0.6639 0.6632 0.6787 0.6682
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large
0.6615 0.6617 0.6619 0.6625 0.6823 0.6638
Qwen2.5-1.5B, Gemma3-1B,
FLAN-T5-Large
0.6575 0.6603 0.6601 0.6603 0.6785 0.6634
Qwen2.5-1.5B, Llama3.2-1B,
mT0-Large
0.6544 0.6513 0.6536 0.6542 0.6732 0.6541
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B, mT0-Large
0.6631 0.6626 0.6642 0.6644 0.6846 0.666
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large, FLAN-T5-Large
0.6582 0.6602 0.6604 0.662 0.6821 0.6618
Llama3.1-8B, Qwen2.5-7B,
FLAN-T5-Large
0.6868 0.6861 0.6883 0.688 0.6982 0.6887
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
0.6885 0.6859 0.6864 0.6865 0.6986 0.686
Llama3.1-8B, Qwen2.5-7B,
Qwen2.5-1.5B
0.6841 0.6858 0.6856 0.6855 0.6964 0.6853
Llama3.2-3B, Qwen2.5-3B,
Gemma3-1B
0.6791 0.6794 0.679 0.6795 0.6936 0.6817
@@ H.4.4 CNN/DailyMail
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.2531 0.2578 0.259 0.8
Qwen2.5-3B 0.252 0.2511 0.2509 0.3
Qwen2.5-1.5B 0.2315 0.2511 0.2545 0.9
Qwen2.5-0.5B 0.2293 0.2416 0.2452 1.0
Gemma3-1B 0.2411 0.2431 0.2441 0.9
Llama3.2-1B 0.2468 0.2479 0.2479 0.5
Llama3.2-3B 0.2505 0.2527 0.2527 0.1
Llama3.1-8B 0.251 0.2532 0.2534 0.9
mT0-Large 0.2513 0.2512 0.2524 0.8
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Llama3.2-1B, mT0-Large 0.2543 0.2494 0.2517 0.1
Llama3.2-1B, Qwen2.5-1.5B 0.2483 0.2457 0.2526 0.3
Qwen2.5-1.5B, Gemma3-1B 0.2454 0.2453 0.2549 1.0
Llama3.2-1B, mT0-Large, Qwen2.5-1.5B 0.2536 0.2448 0.2535 0.4
Qwen2.5-7B, Llama3.1-8B 0.2533 0.2538 0.2566 0.2
Qwen2.5-7B, mT0-Large 0.2533 0.2537 0.2561 0.2
Qwen2.5-7B, Qwen2.5-3B 0.2558 0.2542 0.2568 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Llama3.2-1B, mT0-Large,
Qwen2.5-1.5B
0.2595 0.2611 0.2593 0.2588 0.2537 0.2587
Llama3.2-1B, mT0-Large,
Qwen2.5-0.5B
0.2551 0.2542 0.2533 0.2551 0.2478 0.2523
Qwen2.5-1.5B, Qwen2.5-0.5B,
Gemma3-1B
0.2525 0.2542 0.2482 0.2502 0.251 0.249
Llama3.2-1B, Gemma3-1B,
Qwen2.5-0.5B
0.2504 0.2507 0.2511 0.2519 0.2471 0.2497
Llama3.2-3B, Qwen2.5-3B,
Qwen2.5-0.5B
0.254 0.2563 0.255 0.255 0.2514 0.2539
Qwen2.5-7B, mT0-Large,
Llama3.2-3B
0.2605 0.2598 0.2581 0.2574 0.2556 0.2584
Qwen2.5-7B, Llama3.1-8B,
mT0-Large
0.2574 0.2578 0.2595 0.2585 0.2548 0.2583
Llama3.1-8B, mT0-Large,
Qwen2.5-3B
0.2543 0.2532 0.2543 0.2536 0.2501 0.2535
@@
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.8227 0.818 0.8227 0.0
Qwen2.5-3B 0.7833 0.7853 0.786 0.1
Qwen2.5-1.5B 0.788 0.7813 0.782 0.1
Qwen2.5-0.5B 0.67 0.6593 0.6647 0.0
Gemma3-1B 0.674 0.666 0.6727 0.0
Llama3.2-1B 0.638 0.648 0.6527 0.8
Llama3.2-3B 0.766 0.7687 0.7747 0.8
Llama3.1-8B 0.8207 0.814 0.816 0.1
FLAN-T5-Large 0.7567 0.6453 0.7247 0.0
mT0-Large 0.8267 0.8307 0.8307 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
mT0-Large, Qwen2.5-1.5B 0.8153 0.8207 0.8113 0.0
mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large 0.7807 0.8073 0.8013 0.0
Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B 0.7193 0.724 0.746 0.0
Qwen2.5-7B, Llama3.1-8B 0.8207 0.8213 0.8227 0.0
Qwen2.5-7B, Llama3.1-8B, mT0-Large 0.824 0.8227 0.8233 0.1
Qwen2.5-7B, Llama3.2-3B 0.7867 0.7907 0.788 0.0
Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.7273 0.7167 0.7607 1.0
Qwen2.5-7B, mT0-Large, FLAN-T5-Large 0.7747 0.816 0.8127 0.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
mT0-Large, Qwen2.5-1.5B,
FLAN-T5-Large
0.8313 0.8233 0.8053 0.824 0.8227 0.8247
mT0-Large, Qwen2.5-1.5B,
FLAN-T5-Large, Llama3.2-1B
0.8227 0.826 0.8067 0.8273 0.8167 0.8187
Qwen2.5-1.5B, FLAN-T5-Large,
Gemma3-1B
0.7813 0.79 0.7653 0.7913 0.78 0.786
Qwen2.5-1.5B, FLAN-T5-Large,
Gemma3-1B, Llama3.2-1B
0.7667 0.7893 0.762 0.7887 0.7873 0.7713
Qwen2.5-1.5B, FLAN-T5-Large,
Llama3.2-1B
0.7833 0.7967 0.7573 0.7987 0.782 0.782
Llama3.2-3B, Qwen2.5-3B,
FLAN-T5-Large
0.808 0.808 0.7913 0.8087 0.8187 0.806
Qwen2.5-7B, mT0-Large,
FLAN-T5-Large
0.8427 0.8333 0.8207 0.8347 0.838 0.8433
Qwen2.5-7B, mT0-Large,
Llama3.2-3B
0.8287 0.83 0.8227 0.8287 0.8293 0.832
Qwen2.5-7B, Llama3.1-8B,
mT0-Large
0.836 0.8353 0.8207 0.834 0.8373 0.838
Llama3.1-8B, mT0-Large,
Qwen2.5-1.5B
0.83 0.8313 0.8107 0.832 0.8287 0.83
@@ H.4.6 TriviaQA
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.709 0.7017 0.7073 0.0
Qwen2.5-3B 0.6422 0.6373 0.6373 0.1
Qwen2.5-1.5B 0.5796 0.5752 0.5767 0.1
Qwen2.5-0.5B 0.4149 0.4014 0.4111 0.0
Gemma3-1B 0.4928 0.4915 0.4906 0.1
Llama3.2-1B 0.5249 0.5193 0.5208 0.2
Llama3.2-3B 0.7137 0.7071 0.7113 0.1
Llama3.1-8B 0.7763 0.7711 0.7726 0.1
FLAN-T5-Large 0.4403 0.435 0.4438 0.0
mT0-Large 0.3637 0.3623 0.3676 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-1.5B, Llama3.2-1B 0.5641 0.5708 0.5699 0.1
Llama3.2-1B, Gemma3-1B 0.5193 0.5229 0.5239 0.2
Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B 0.5519 0.5559 0.5571 0.0
Qwen2.5-1.5B, Llama3.2-1B, FLAN-T5-Large 0.5322 0.547 0.5487 0.0
Llama3.1-8B, Qwen2.5-7B 0.7301 0.7266 0.7401 0.0
Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.5807 0.5478 0.6098 0.0
Qwen2.5-7B, Llama3.2-3B 0.7154 0.7216 0.7203 0.0
Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.6377 0.6207 0.6465 0.0
Qwen2.5-7B, Gemma3-1B 0.5925 0.5922 0.6004 0.0
Llama3.2-3B, Qwen2.5-3B 0.6857 0.6852 0.6883 0.1
Llama3.1-8B, Llama3.2-3B, Qwen2.5-3B 0.7026 0.6889 0.7158 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Qwen2.5-1.5B, Llama3.2-1B,
Gemma3-1B
0.5314 0.5427 0.4969 0.5429 0.533 0.5307
Qwen2.5-1.5B, Llama3.2-1B,
FLAN-T5-Large
0.5198 0.5319 0.4944 0.5322 0.5191 0.5208
Llama3.2-1B, Gemma3-1B,
FLAN-T5-Large
0.4872 0.4968 0.4604 0.4973 0.4891 0.4857
Qwen2.5-1.5B, Llama3.2-1B,
Gemma3-1B, FLAN-T5-Large
0.5303 0.541 0.4925 0.541 0.5332 0.5231
Llama3.1-8B, Qwen2.5-7B,
Llama3.2-3B
0.743 0.7494 0.6964 0.7496 0.7407 0.7227
Qwen2.5-7B, Llama3.2-3B,
Gemma3-1B
0.6698 0.68 0.6143 0.6793 0.6708 0.6636
Qwen2.5-7B, Llama3.2-3B,
Qwen2.5-3B
0.6833 0.7007 0.6154 0.6997 0.695 0.6822
Llama3.1-8B, Llama3.2-3B,
Qwen2.5-3B
0.7391 0.7452 0.7009 0.745 0.7341 0.7181
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
0.711 0.7158 0.6969 0.7164 0.6803 0.668

### v3

@@ H.4 Performances at 40% Target Model FLOPs
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.7337 0.734 0.7336 0.7341 0.7382 0.7343 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.7345 0.7351 0.734 0.735 0.7402 0.7353 Qwen2.5-1.5B, Gemma3-1B, FLAN-T5-Large 0.7366 0.7366 0.7359 0.737 0.7415 0.7376 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.731 0.7323 0.7312 0.7328 0.7368 0.7335 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.7337 0.734 0.7338 0.7342 0.7398 0.7347 Gemma3-1B, Qwen2.5-1.5B, mT0-Large, FLAN-T5-Large 0.735 0.7361 0.734 0.7358 0.741 0.7359 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.7443 0.7443 0.7431 0.744 0.7473 0.7449 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.7429 0.7431 0.7421 0.7432 0.7466 0.7446 Llama3.1-8B, Qwen2.5-7B, Qwen2.5-1.5B 0.7438 0.7437 0.7443 0.744 0.7455 0.7444 Llama3.2-3B, Qwen2.5-3B, Gemma3-1B 0.7379 0.7386 0.7377 0.7382 0.7423 0.7398
@@
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.684 0.6903 0.6887 0.1 Qwen2.5-3B 0.665 0.6745 0.673 0.2 Qwen2.5-1.5B 0.6579 0.6659 0.6667 0.3 Qwen2.5-0.5B 0.5947 0.6027 0.6014 0.3 Gemma3-1B 0.6664 0.6701 0.67 0.2 Llama3.2-1B 0.6467 0.6528 0.6522 0.4 Llama3.2-3B 0.6774 0.6872 0.6843 0.1 Llama3.1-8B 0.6909 0.6956 0.6968 0.6 FLAN-T5-Large 0.5834 0.5896 0.5894 0.4 mT0-Large 0.6107 0.6186 0.6154 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Gemma3-1B, Qwen2.5-1.5B 0.6673 0.6691 0.6705 0.1 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6657 0.6692 0.6692 0.1 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6625 0.6649 0.6663 0.0 Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.6645 0.6626 0.667 0.0 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6592 0.6634 0.6638 0.1 Llama3.1-8B, Qwen2.5-7B 0.6869 0.6903 0.6898 0.1 Llama3.1-8B, Qwen2.5-3B 0.6818 0.6854 0.6876 0.2 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6855 0.689 0.6892 0.1 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6843 0.6883 0.6899 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6623 0.6622 0.6639 0.6632 0.6787 0.6682 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6615 0.6617 0.6619 0.6625 0.6823 0.6638 Qwen2.5-1.5B, Gemma3-1B, FLAN-T5-Large 0.6575 0.6603 0.6601 0.6603 0.6785 0.6634 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6544 0.6513 0.6536 0.6542 0.6732 0.6541 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6631 0.6626 0.6642 0.6644 0.6846 0.666 Gemma3-1B, Qwen2.5-1.5B, mT0-Large, FLAN-T5-Large 0.6582 0.6602 0.6604 0.662 0.6821 0.6618 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6868 0.6861 0.6883 0.688 0.6982 0.6887 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6885 0.6859 0.6864 0.6865 0.6986 0.686 Llama3.1-8B, Qwen2.5-7B, Qwen2.5-1.5B 0.6841 0.6858 0.6856 0.6855 0.6964 0.6853 Llama3.2-3B, Qwen2.5-3B, Gemma3-1B 0.6791 0.6794 0.679 0.6795 0.6936 0.6817
@@ H.4.4 CNN/DailyMail
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.2531 0.2578 0.259 0.8 Qwen2.5-3B 0.252 0.2511 0.2509 0.3 Qwen2.5-1.5B 0.2315 0.2511 0.2545 0.9 Qwen2.5-0.5B 0.2293 0.2416 0.2452 1.0 Gemma3-1B 0.2411 0.2431 0.2441 0.9 Llama3.2-1B 0.2468 0.2479 0.2479 0.5 Llama3.2-3B 0.2505 0.2527 0.2527 0.1 Llama3.1-8B 0.251 0.2532 0.2534 0.9 mT0-Large 0.2513 0.2512 0.2524 0.8
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Llama3.2-1B, mT0-Large 0.2543 0.2494 0.2517 0.1 Llama3.2-1B, Qwen2.5-1.5B 0.2483 0.2457 0.2526 0.3 Qwen2.5-1.5B, Gemma3-1B 0.2454 0.2453 0.2549 1.0 Llama3.2-1B, mT0-Large, Qwen2.5-1.5B 0.2536 0.2448 0.2535 0.4 Qwen2.5-7B, Llama3.1-8B 0.2533 0.2538 0.2566 0.2 Qwen2.5-7B, mT0-Large 0.2533 0.2537 0.2561 0.2 Qwen2.5-7B, Qwen2.5-3B 0.2558 0.2542 0.2568 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Llama3.2-1B, mT0-Large, Qwen2.5-1.5B 0.2595 0.2611 0.2593 0.2588 0.2537 0.2587 Llama3.2-1B, mT0-Large, Qwen2.5-0.5B 0.2551 0.2542 0.2533 0.2551 0.2478 0.2523 Qwen2.5-1.5B, Qwen2.5-0.5B, Gemma3-1B 0.2525 0.2542 0.2482 0.2502 0.251 0.249 Llama3.2-1B, Gemma3-1B, Qwen2.5-0.5B 0.2504 0.2507 0.2511 0.2519 0.2471 0.2497 Llama3.2-3B, Qwen2.5-3B, Qwen2.5-0.5B 0.254 0.2563 0.255 0.255 0.2514 0.2539 Qwen2.5-7B, mT0-Large, Llama3.2-3B 0.2605 0.2598 0.2581 0.2574 0.2556 0.2584 Qwen2.5-7B, Llama3.1-8B, mT0-Large 0.2574 0.2578 0.2595 0.2585 0.2548 0.2583 Llama3.1-8B, mT0-Large, Qwen2.5-3B 0.2543 0.2532 0.2543 0.2536 0.2501 0.2535
@@
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.8227 0.818 0.8227 0.0 Qwen2.5-3B 0.7833 0.7853 0.786 0.1 Qwen2.5-1.5B 0.788 0.7813 0.782 0.1 Qwen2.5-0.5B 0.67 0.6593 0.6647 0.0 Gemma3-1B 0.674 0.666 0.6727 0.0 Llama3.2-1B 0.638 0.648 0.6527 0.8 Llama3.2-3B 0.766 0.7687 0.7747 0.8 Llama3.1-8B 0.8207 0.814 0.816 0.1 FLAN-T5-Large 0.7567 0.6453 0.7247 0.0 mT0-Large 0.8267 0.8307 0.8307 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile mT0-Large, Qwen2.5-1.5B 0.8153 0.8207 0.8113 0.0 mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large 0.7807 0.8073 0.8013 0.0 Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B 0.7193 0.724 0.746 0.0 Qwen2.5-7B, Llama3.1-8B 0.8207 0.8213 0.8227 0.0 Qwen2.5-7B, Llama3.1-8B, mT0-Large 0.824 0.8227 0.8233 0.1 Qwen2.5-7B, Llama3.2-3B 0.7867 0.7907 0.788 0.0 Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.7273 0.7167 0.7607 1.0 Qwen2.5-7B, mT0-Large, FLAN-T5-Large 0.7747 0.816 0.8127 0.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large 0.8313 0.8233 0.8053 0.824 0.8227 0.8247 mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large, Llama3.2-1B 0.8227 0.826 0.8067 0.8273 0.8167 0.8187 Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B 0.7813 0.79 0.7653 0.7913 0.78 0.786 Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B, Llama3.2-1B 0.7667 0.7893 0.762 0.7887 0.7873 0.7713 Qwen2.5-1.5B, FLAN-T5-Large, Llama3.2-1B 0.7833 0.7967 0.7573 0.7987 0.782 0.782 Llama3.2-3B, Qwen2.5-3B, FLAN-T5-Large 0.808 0.808 0.7913 0.8087 0.8187 0.806 Qwen2.5-7B, mT0-Large, FLAN-T5-Large 0.8427 0.8333 0.8207 0.8347 0.838 0.8433 Qwen2.5-7B, mT0-Large, Llama3.2-3B 0.8287 0.83 0.8227 0.8287 0.8293 0.832 Qwen2.5-7B, Llama3.1-8B, mT0-Large 0.836 0.8353 0.8207 0.834 0.8373 0.838 Llama3.1-8B, mT0-Large, Qwen2.5-1.5B 0.83 0.8313 0.8107 0.832 0.8287 0.83
@@ H.4.6 TriviaQA
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.709 0.7017 0.7073 0.0 Qwen2.5-3B 0.6422 0.6373 0.6373 0.1 Qwen2.5-1.5B 0.5796 0.5752 0.5767 0.1 Qwen2.5-0.5B 0.4149 0.4014 0.4111 0.0 Gemma3-1B 0.4928 0.4915 0.4906 0.1 Llama3.2-1B 0.5249 0.5193 0.5208 0.2 Llama3.2-3B 0.7137 0.7071 0.7113 0.1 Llama3.1-8B 0.7763 0.7711 0.7726 0.1 FLAN-T5-Large 0.4403 0.435 0.4438 0.0 mT0-Large 0.3637 0.3623 0.3676 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-1.5B, Llama3.2-1B 0.5641 0.5708 0.5699 0.1 Llama3.2-1B, Gemma3-1B 0.5193 0.5229 0.5239 0.2 Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B 0.5519 0.5559 0.5571 0.0 Qwen2.5-1.5B, Llama3.2-1B, FLAN-T5-Large 0.5322 0.547 0.5487 0.0 Llama3.1-8B, Qwen2.5-7B 0.7301 0.7266 0.7401 0.0 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.5807 0.5478 0.6098 0.0 Qwen2.5-7B, Llama3.2-3B 0.7154 0.7216 0.7203 0.0 Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.6377 0.6207 0.6465 0.0 Qwen2.5-7B, Gemma3-1B 0.5925 0.5922 0.6004 0.0 Llama3.2-3B, Qwen2.5-3B 0.6857 0.6852 0.6883 0.1 Llama3.1-8B, Llama3.2-3B, Qwen2.5-3B 0.7026 0.6889 0.7158 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B 0.5314 0.5427 0.4969 0.5429 0.533 0.5307 Qwen2.5-1.5B, Llama3.2-1B, FLAN-T5-Large 0.5198 0.5319 0.4944 0.5322 0.5191 0.5208 Llama3.2-1B, Gemma3-1B, FLAN-T5-Large 0.4872 0.4968 0.4604 0.4973 0.4891 0.4857 Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B, FLAN-T5-Large 0.5303 0.541 0.4925 0.541 0.5332 0.5231 Llama3.1-8B, Qwen2.5-7B, Llama3.2-3B 0.743 0.7494 0.6964 0.7496 0.7407 0.7227 Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 0.6698 0.68 0.6143 0.6793 0.6708 0.6636 Qwen2.5-7B, Llama3.2-3B, Qwen2.5-3B 0.6833 0.7007 0.6154 0.6997 0.695 0.6822 Llama3.1-8B, Llama3.2-3B, Qwen2.5-3B 0.7391 0.7452 0.7009 0.745 0.7341 0.7181 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.711 0.7158 0.6969 0.7164 0.6803 0.668